# Quantitative Introspection in Language Models: Tracking Internal States Across Conversation

### 19 Mar 2026 v1

@@ 4.3 Introspection is present from the first turn and evolves over time
Figs. 3B–C show the turn-wise introspective strength (Spearman $\rho$) and introspective fidelity (isotonic $R^{2}$) for each concept. Introspection is already present at turn 1: wellbeing, interest, focus, and impulsivity all show positive turn-1 correlations ($\rho=0.52$, $0.55$, $0.26$, and $0.65$, respectively), and three are already significant at turn 1 (wellbeing $p=5.46\times 10^{-4}$; interest $p=2.37\times 10^{-4}$; impulsivity $p=6.80\times 10^{-6}$; focus $p=0.11$). These three concepts show significant introspection across all ten turns (all per-turn $p\leq 0.019$; all survive within-concept BH correction), suggesting this capacity is robust over time. The focus direction does not show significant introspection at turn 1, but becomes significant in 8/10 turns nominally and 6/10 turns after within-concept BH correction (significant-turn $\rho$ range: $0.32$ to $0.53$). Together, these findings show that the observed correlation is not explained by joint temporal drift alone, but also appears at the level of individual turns throughout conversation. For some concepts, it appears from the beginning of a conversation, without needing multi-turn context to build up the association.
Since we now have the trajectory of introspective capacity over time for each concept, an interesting follow-up question is whether this effect changes over time or is otherwise stable. Strikingly, we found that introspection does vary significantly as the conversation progresses, but this evolution is concept-dependent. Wellbeing, interest, and focus show an overall increase in introspective fidelity from early to late turns (Fig. 3C, $\Delta R^{2}=+0.31$, +0.27+0.27, and +0.17+0.17 from turn 1 to turn 10 respectively), while impulsivity shows the opposite pattern ($\Delta R^{2}=-0.28$), with self-report fidelity weakening over time. Mixed-effects interaction models (logit_report $\sim$ probe $\times$ turn ++ (1||conversation)) confirm that the probe-report coupling changes significantly over time for all four concepts (interaction $p<0.01$ in all cases).
Related prior work has studied temporal changes in factual confidence or diachronic self-consistency (Zhang et al., 2026; Prestes, 2025), but those settings concern either correctness with respect to the external world or stability of self-description across prompts. To the best of our knowledge, this is the first report of emotive introspective capacity changing over conversational time in LLMs.
4.4 Self-steering confirms causal dependence
@@
Fig. 3D presents the self-steering results: mean logit-based self-report as a function of steering strength $\alpha$, for each concept steered along the corresponding probe’s direction. In all four cases, self-report increases monotonically with $\alpha$ (LMM alpha slopes: wellbeing =0.19=0.19, interest =0.25=0.25, focus =0.40=0.40, impulsivity =0.067=0.067; all $p<10^{-12}$), confirming that shifting the internal state along the probe-defined direction shifts the model’s report in the predicted direction. Focus, interest and wellbeing show the largest absolute shifts (between 1.5 and 3.2 rating points across the full alpha range), while impulsivity shows a smaller but reliable effect. In all four cases, self-steering has an effect in both directions: steering the model towards the negative pole of the probe decreases the self-reported score, while steering the model towards the positive pole of the probe increases the self-reported score.
Having shown that self-steering changes a state’s self-report as predicted, we next asked whether this intervention also modulates how that report evolves over time. Fig. 3E shows the drift magnitude (last-turn minus first-turn self-report) as a function of $\alpha$. The relationship is concept-specific: the absolute magnitude of drift increases with $\alpha$ for wellbeing and impulsivity (LMM slope =0.012=0.012, $p<10^{-5}$; LMM slope =0.028=-0.028, $p<10^{-52}$), while interest and focus drift decrease with $\alpha$ (per-conversation slope means =0.095=-0.095 and 0.11-0.11, both $p<10^{-13}$; fallback tests are applied because the LMMs are singular, see Methods). The per-concept steering-drift curves are shown in Appendix D. Together, these results confirm that the coupling between self-report and internal state is causal: moving the probe-defined direction shifts the corresponding report as expected, and the relationship holds across the full range of steering strengths. A causal relation between internal state and self-report supports the claim that the model is performing introspection on these four concepts.
4.5 Cross-concept steering can improve introspection
@@
Having shown that the model possesses a basal capacity for introspection, we next asked whether this ability is optimal, or if it can be improved. Introspective fidelity varies across concepts, and sometimes it can be relatively low (e.g., $R^{2}=0.12$ for focus vs. $0.54$ for interest). We reasoned that it could be possible to intervene the model by steering it towards some direction in latent space so that the association between self-reports and the underlying emotive states is improved. However, it could also be possible that basal introspective fidelity was optimal, so steering the model away from its default state would only degrade this ability. Our goal here was to show that improvement is possible for at least some concepts by using a steering intervention, not to find an optimal steering direction that would improve introspection for all concepts. Taking advantage of the fact that we already have four interpretable internal directions that we have validated, we decided to perform cross-steering experiments, where we steer the model along one of the four concept directions while measuring introspection for another. This yields a $4\times 4$ matrix of steering-concept $\times$ measured-concept conditions (16 cells, each evaluated at five alpha values), for a total of 80 experimental conditions. For each condition, we recorded the isotonic $R^{2}$ between logit-based self-report and previous turn probe score for the measured concept. Then we evaluated if this metric showed significant monotonic variation when we varied alpha along the steering concept direction (i.e. if introspective fidelity could be modulated reliably by steering the model along a linear direction in latent space).
@@
It is important to note that introspection improvement appears to be concept-pair-specific: in most cells of the $4\times 4$ matrix, steering has no significant effect over introspection fidelity. However, the fact that this effect exists at any single one of these conditions is already a significant finding: since the steering effect is constant for a given experiment, an increased association between self-reports and corresponding internal states cannot be explained away as an artifact of intervention, but reflects an improvement in introspective fidelity of that emotive state. However, this does suggest that introspection is not a single ability that can be modulated equally for all internal directions, but might depend on the internal geometry relating the steered direction to the measured concept.
@@
We next wanted to better understand what drives the improvement in the two selected conditions. Conceptually, one could decompose introspection into two components: (i) the informativeness of the internal state itself (a noisier or flatter representation provides less signal to report on), and (ii) the quality of the report given that state (a model with rich internal variation might still produce collapsed or uninformative reports). To distinguish these, we attempted to examine how steering affects each component separately. For this analysis, we focus on Shannon entropy of the previous-turn probe scores (Fig. 4F) and the logit-based self-reports (Fig. 4G), as a measure of how much usable variation each channel contains.
@@ 4.6 Generalization across model sizes and families
Figure 5: Results generalize unevenly across model scales and families. Panel A shows isotonic $R^{2}$ versus model size; panel B shows Spearman $\rho$ versus model size. Introspection increases strongly with size for wellbeing and interest, but remains weak for focus and impulsivity. Panel C shows probe score versus logit self-report for the two strongest LLaMA 8B probes, wellbeing and interest, with black isotonic fits. Both show near-ceiling introspection ($\rho=0.93$ and $0.96$; isotonic $R^{2}=0.90$ and $0.93$), and mixed-effects probe slopes are strongly positive in both cases ($p<10^{-10}$). Panel D shows the pooled Spearman $\rho(\alpha,\text{self-report})$ from the self-steering analysis across concept and LLaMA size, as a descriptive sign-validation heatmap. Panel E shows mean isotonic $R^{2}$ across the subset of concept-model pairs that pass the steering-sign validation filter, with bootstrap 95% CIs. Mean validated $R^{2}$ increases from 1B to 3B to 8B (0.12, 0.37, 0.61); a pooled LMM over validated cells with probe-$z\times\log(\text{size})$, concept intercepts and concept-specific probe slopes, and random intercept by conversation is strongly positive ($\beta=0.29$, $p=5.55\times 10^{-99}$). Panels F and G show probe-score drift across turns for the wellbeing direction and the corresponding mean first-to-last probe drift by size; probe-drift magnitude increases with scale (bootstrap slope versus $\log(\text{size})=0.041$, $p<2\times 10^{-4}$). Panels H and I show the analogous logit self-report drift analyses: mixed-effects turn slopes are positive for all three sizes (1B: $0.159$, $p=1.13\times 10^{-56}$; 3B: $0.038$, $p=4.01\times 10^{-30}$; 8B: $0.141$, $p=8.85\times 10^{-128}$), but unlike probe drift, report-drift magnitude does not increase with scale, and the size-slope is negative (mean slope =0.23=-0.23, $p=0.023$). Panel J shows logit self-report drift through turns for wellbeing in Gemma and Qwen; mixed-effects turn slopes are positive in both families (0.056 and 0.026, both $p<10^{-3}$). Panels K and L show probe score versus logit self-report for Gemma and Qwen, respectively. Qwen shows stronger introspection than Gemma ($\rho=0.49$ vs. $0.28$; isotonic $R^{2}=0.76$ vs. $0.11$), and mixed-effects probe slopes are positive in both cases ($p=1.19\times 10^{-84}$ and $1.33\times 10^{-13}$, respectively). Panels M and N show turn-wise isotonic $R^{2}$ and turn-wise Spearman $\rho$ for the same cross-family comparison; Qwen shows a strong first-to-last decline in turn-wise isotonic $R^{2}$ ($\Delta=-0.44$, cluster-bootstrap $p=0.001$), whereas the corresponding first-to-last $\rho$ changes are not significant in either family.
@@
One plausible explanation for these results is that probe quality is degraded for the 1B and 8B models. To be able to ensure fair comparison between different models, we trained same-concept probes with identical system prompts, questions, and evaluation sentences for all models. However, these were first finetuned for the 3B model, as this was the main model used throughout this work, and it’s possible that equivalent probe training conditions are less effective in other models. To account for this, we again performed self-steering experiments for all concepts and models, to see if steering a model towards an interpretable pole would change its self-report in the predicted direction reliably (see Appendix F for details). This was the case for all models in the two concepts where we saw consistent improvement with scale (Fig. 5D, note positive $\rho$ values between self-report and steering alpha; wellbeing: all LMM slope tests $p<10^{-16}$; interest: $p=7.56\times 10^{-9}$ for 1B, and $p<10^{-16}$ for 3B and 8B). However, steering in the direction of the focus probe for the 1B model and the impulsivity probe for the 8B model caused an inverse change in self-reported rating (i.e. when steering towards the positive probe pole, the rating became more negative, and vice versa; note negative $\rho$ values between self-report and steering alpha in Fig. 5D). This could be a sign that these probes are low-quality, or that these models are not able to report on these concepts reliably. In any case, we decided to exclude these two cases, retaining only concept-model pairs where self-steering shifts the report in the expected direction. Fig. 5E shows mean introspective fidelity across validated concept pairs at each model size: 0.12 (1B), 0.37 (3B), and 0.61 (8B) (pooled LMM over validated cells with probe-$z\times\log(\text{size})$, controlling for concept and conversation: $\beta=0.29$, $p=5.55\times 10^{-99}$). Overall, this shows that introspective ability can scale with model size, although this was particularly evident for only two out of four concepts tested.
@@ 5.1 Scope of the claim
Our central claim is deliberately limited. We do not claim that these models have conscious felt experience, nor that a numeric self-report gives direct access to anything like human phenomenology. Instead, we show that some instruction-tuned LLMs contain measurable internal representations along emotive concept directions, and that these representations can be meaningfully queried through self-report in a way that is both quantitatively coupled to probe-defined state and causally dependent on that interpertable internal direction. This bounded framing follows the criterion that introspection should involve causal dependence between an internal state and the report about that state, rather than mere production of introspective-sounding language (Comsa and Shanahan, 2025), and it remains agnostic about whether such abilities imply anything like conscious experience (McClelland, 2024).
@@ 5.2 Numeric introspection as a complement to white-box methods
To make these contributions more useful for future research, we share an open-source library that allows this framework to be used with any concept, including methods for probe training, multi-probe recording, steering and behavioral testing; we also propose a method for overcoming the natively collapsed ratings output by models when asked to self-report emotive states, making these reports more informative and introspection work a lot better as a tool.
@@ 5.3 Introspection is concept-specific and only locally improvable in our experiments
The cross-steering results are therefore best read as an existence proof, not as evidence for a single, general “introspection knob”. We show that modest representational interventions can in some cases improve introspective fidelity in targeted concept pairs, which is already important: it demonstrates that basal introspective performance is not always maximal, and that some failure cases may be solvable by intervention rather than showing a complete absence of introspective capacity. But we did not search systematically for a universal steering direction that improves introspection across concepts. In pilot analyses we also examined additional interpretable directions, including truthfulness- and authenticity-style directions, but did not observe robust cross-concept introspection gains there (results not shown). The current evidence therefore points toward local, pair-specific improvement rather than a single globally tunable faculty. However, future work should still treat the search of this global faculty as an open problem.
@@ 5.5 Generalization to other models and limitations of this work
Even under that constraint, we find that introspective capacity generalizes to different model sizes and families. First, we find that introspective capacity grows with scale in two out of the four concepts we tested, reaching near-optimal introspection in the 8B LLaMA model (Fig. 5). Second, we replicate the core results outside the LLaMA family using the wellbeing concept in Qwen, while Gemma shows weaker but still detectable evidence. Therefore, we claim that introspective ability is not confined to a single architecture, at least for the concepts tested here. It is also important to note that we detected introspective ability in relatively small models. Much of the strongest prior evidence for mechanistically grounded introspection has involved larger or more heavily scaffolded systems, whereas our main experiments center on a 3B model and still recover causal and temporal structure (Li et al., 2025; Rivera and Africa, 2025; Lindsey, 2026).
@@
In our experiments, the emotive internal states that we study are partly determined by the previous context of the conversation. One could therefore argue that the models could be merely tracking and reporting features of the dialogue rather than performing true introspection, which requires privileged access. Our self-steering experiments show that self-reports change when the corresponding internal state changes, even when conversational context is the same. This shows that some degree of privileged access is necessary to explain the model’s output. However, part of the information the model is using to give self-reported ratings could be coming purely from conversational context. Prior work has shown that models can perform introspection in settings that do not depend on any specific natural conversation history (Lindsey, 2026; Pearson-Vogel et al., 2026; Li et al., 2025), which supports the claim that reports measured here are truly introspective. But it’s important to note that, to test introspective capacity of models in the contexts where they will be used (e.g. during conversations and over time), it is necessary to study internal states that are related to that context, as we do in this work.

### 11 Apr 2026 v2

@@ 4.2.1 Introspection is present from the first turn and evolves over time
Fig. 3B shows the turn-wise introspective fidelity (solid lines; isotonic $R^{2}$) and introspective strength (dashed lines; Spearman $\rho$) for each concept (see Section 3.5). Introspection is already present at turn 1: wellbeing, interest, focus, and impulsivity all show positive turn-1 correlations ($\rho=0.52$, $0.55$, $0.26$, and $0.65$, respectively), and three are already significant at turn 1 (wellbeing $p=5.46\times 10^{-4}$; interest $p=2.37\times 10^{-4}$; impulsivity $p=6.80\times 10^{-6}$; focus $p=0.11$). These three concepts show significant introspection across all ten turns (all per-turn $p\leq 0.019$; all survive within-concept BH correction), suggesting this capacity is robust over time. The focus direction does not show significant introspection at turn 1, but becomes significant in 8/10 turns nominally and 6/10 turns after within-concept BH correction (significant-turn $\rho$ range: $0.32$ to $0.53$). Together, these findings show that the observed correlation is not explained by joint temporal drift alone, but also appears at the level of individual turns throughout conversation. For some concepts, it appears from the beginning of a conversation, without needing multi-turn context to build up the association.
Since we now have the trajectory of introspective capacity over time for each concept, an interesting follow-up question is whether this effect changes over time or is otherwise stable. Strikingly, we found that introspection does vary significantly as the conversation progresses, but this evolution is concept-dependent. Wellbeing, interest, and focus show an overall increase in introspective fidelity from early to late turns (Fig. 3B, $\Delta R^{2}=+0.31$, +0.27+0.27, and +0.17+0.17 from turn 1 to turn 10 respectively), while impulsivity shows the opposite pattern ($\Delta R^{2}=-0.28$), with self-report fidelity weakening over time. Mixed-effects interaction models (logit_report $\sim$ probe $\times$ turn ++ (1||conversation)) confirm that the probe-report coupling changes significantly over time for all four concepts (interaction $p<0.01$ in all cases).
Related prior work has studied temporal changes in factual confidence or diachronic self-consistency (Zhang et al., 2026; Prestes, 2025), but those settings concern either correctness with respect to the external world or stability of self-description across prompts. To the best of our knowledge, this is the first report of emotive introspective capacity changing over conversational time in LLMs. The pooled introspection results discussed above are also reported in full in Appendix G, Table 2.
4.3 Steering reveals causal introspection
@@
Fig. 4A presents the same-concept steering results: mean logit-based self-report as a function of steering strength $\alpha$, for each concept steered along the corresponding probe’s direction. In all four cases, self-report increases monotonically with $\alpha$ (LMM alpha slopes: wellbeing =0.19=0.19, interest =0.25=0.25, focus =0.40=0.40, impulsivity =0.067=0.067; all $p<10^{-12}$), confirming that shifting the internal state along the probe-defined direction shifts the model’s report in the predicted direction. Focus, interest and wellbeing show the largest absolute shifts (between 1.5 and 3.2 rating points across the full alpha range), while impulsivity shows a smaller but reliable effect. In all four cases, same-concept steering has an effect in both directions: steering the model towards the negative pole of the probe decreases the self-reported score, while steering the model towards the positive pole of the probe increases the self-reported score. Together, these results confirm that the coupling between self-report and internal state is causal: moving the probe-defined direction shifts the corresponding report as expected, and the relationship holds across the full range of steering strengths. A causal relation between internal state and self-report supports the claim that the model is performing introspection on these four concepts.
Cross-concept steering can improve introspection
@@
Having shown that the model possesses a basal capacity for introspection, we next asked whether this ability is optimal, or if it can be improved. Introspective fidelity varies across concepts, and sometimes it can be relatively low (e.g., $R^{2}=0.12$ for focus vs. $0.54$ for interest). If steering the model in any direction, away from its default state, can improve that coupling, then basal introspection is not optimal. Because arbitrary directions are harder to interpret, we used the four validated concept directions as steering axes. Our goal here was to show that improvement is possible for at least some concepts by using a steering intervention, not to find an optimal steering direction that would improve introspection for all concepts. We therefore performed cross-concept steering experiments, where we steer the model along one of the four concept directions while measuring introspection for another. This yields a $4\times 4$ matrix of steering-concept $\times$ measured-concept conditions (16 cells, each evaluated at five alpha values), for a total of 80 experimental conditions. For each condition, we recorded introspective fidelity (isotonic $R^{2}$ between logit-based self-report and previous turn probe score) for the measured concept. Then we evaluated if this metric showed significant monotonic variation when we varied alpha along the steering concept direction (i.e. if introspective fidelity could be modulated reliably by steering the model along a linear direction in latent space).
@@
It is important to note that introspection improvement appears to be concept-pair-specific: in most cells of the $4\times 4$ matrix, steering has no significant effect on introspective fidelity. However, the fact that this effect exists at any single one of these conditions is already a significant finding: since the steering effect is constant for a given experiment, an increased association between self-reports and corresponding internal states cannot be explained away as an artifact of intervention, but reflects an improvement in introspective fidelity of that emotive state. However, this does suggest that introspection is not a single ability that can be modulated equally for all internal directions, but might depend on the internal geometry relating the steered direction to the measured concept.
@@
We next wanted to better understand what drives the improvement in the two selected conditions. Conceptually, one could decompose introspection into two components: (i) the informativeness of the internal state itself (a noisier or flatter representation provides less signal to report on), and (ii) the quality of the report given that state (a model with rich internal variation might still produce collapsed or uninformative reports). To distinguish these, we attempted to examine how steering affects each component separately. For this analysis, we focus on Shannon entropy of the previous-turn probe scores and the logit-based self-reports (Fig. 4F), as a measure of how much usable variation each channel contains.
@@ 4.4 Generalization across model sizes and families
Figure 5: Results generalize unevenly across model scales and families. Panels A–B show isotonic $R^{2}$ and Spearman $\rho$ versus model size, computed from the $\alpha=0$ slice of the same-concept steering runs to keep the protocol matched across LLaMA 1B, 3B, and 8B. Introspection increases strongly with size for wellbeing and interest, but remains weak for focus and impulsivity. Panel C shows probe score versus logit self-report for the two strongest LLaMA 8B probes, wellbeing and interest, with black isotonic fits. Both show near-ceiling introspection ($\rho=0.93$ and $0.96$; isotonic $R^{2}=0.90$ and $0.93$), and mixed-effects probe slopes are strongly positive in both cases ($p<10^{-10}$). Panel D shows the pooled Spearman $\rho(\alpha,\text{self-report})$ from the same-concept steering analysis across concept and LLaMA size as a descriptive sign-validation heatmap. Panel E shows mean isotonic $R^{2}$ across the subset of concept-model pairs that pass the steering-sign validation filter, with bootstrap 95% CIs. Mean validated $R^{2}$ increases from 1B to 3B to 8B (0.12, 0.37, 0.61); a pooled LMM over validated cells with probe-$z\times\log(\text{size})$, concept intercepts and concept-specific probe slopes, and random intercept by conversation is strongly positive ($\beta=0.29$, $p=5.55\times 10^{-99}$). Panels F and G show probe-score drift across turns for the wellbeing direction and the corresponding mean first-to-last probe drift by size; probe-drift magnitude increases with scale (bootstrap slope versus $\log(\text{size})=0.041$, $p<2\times 10^{-4}$). Panels H and I show the analogous logit self-report drift analyses: mixed-effects turn slopes are positive for all three sizes (1B: $0.159$, $p=1.13\times 10^{-56}$; 3B: $0.038$, $p=4.01\times 10^{-30}$; 8B: $0.141$, $p=8.85\times 10^{-128}$), but unlike probe drift, report-drift magnitude does not increase with scale, and the size-slope is negative (mean slope =0.23=-0.23, $p=0.023$). Panel J shows logit self-report drift through turns for wellbeing in Gemma and Qwen; mixed-effects turn slopes are positive in both families (0.056 and 0.026, both $p<10^{-3}$). Panels K and L show probe score versus logit self-report for Gemma and Qwen. Qwen shows stronger introspection than Gemma ($\rho=0.49$ vs. $0.28$; isotonic $R^{2}=0.76$ vs. $0.11$), and mixed-effects probe slopes are positive in both cases ($p=1.19\times 10^{-84}$ and $1.33\times 10^{-13}$). Panels M and N show turn-wise isotonic $R^{2}$ and turn-wise Spearman $\rho$ for the same cross-family comparison; Qwen shows a strong first-to-last decline in turn-wise isotonic $R^{2}$ ($\Delta=-0.44$, cluster-bootstrap $p=0.001$), whereas the corresponding first-to-last $\rho$ changes are not significant in either family.
@@
One plausible explanation for these results is that probe quality is degraded for the 1B and 8B models. To be able to ensure fair comparison between different models, we trained same-concept probes with identical system prompts, questions, and evaluation sentences for all models. However, these were first fine-tuned for the 3B model, as this was the main model used throughout this work, and it’s possible that equivalent probe training conditions are less effective in other models. To account for this, we again performed same-concept steering experiments for all concepts and models, to see if steering a model towards an interpretable pole would change its self-report in the predicted direction reliably (see Appendix F for details). This was the case for all models in the two concepts where we saw consistent improvement with scale (Fig. 5D, note positive $\rho$ values between self-report and steering alpha; wellbeing: all LMM slope tests $p<10^{-16}$; interest: $p=7.56\times 10^{-9}$ for 1B, and $p<10^{-16}$ for 3B and 8B). However, steering in the direction of the focus probe for the 1B model and the impulsivity probe for the 8B model caused an inverse change in self-reported rating (i.e. when steering towards the positive probe pole, the rating became more negative, and vice versa; note negative $\rho$ values between self-report and steering alpha in Fig. 5D). This could be a sign that these probes are low-quality, or that these models are not able to report on these concepts reliably. In any case, we decided to exclude these two cases, retaining only concept-model pairs where same-concept steering shifts the report in the expected direction. Fig. 5E shows mean introspective fidelity across validated concept pairs at each model size: 0.12 (1B), 0.37 (3B), and 0.61 (8B) (pooled LMM over validated cells with probe-$z\times\log(\text{size})$, controlling for concept and conversation: $\beta=0.29$, $p=5.55\times 10^{-99}$). Overall, this shows that introspective ability can scale with model size, although this was particularly evident for only two out of four concepts tested.
@@ 5.1 Scope of the claim
Our central claim is deliberately limited. We do not claim that these models have conscious felt experience, nor that a numeric self-report gives direct access to anything like human phenomenology. Instead, we show that some instruction-tuned LLMs contain measurable internal representations along emotive concept directions, and that these representations can be meaningfully queried through self-report in a way that is both quantitatively coupled to probe-defined state and causally dependent on that interpretable internal direction. This bounded framing follows the criterion that introspection should involve causal dependence between an internal state and the report about that state, rather than mere production of introspective-sounding language (Comsa and Shanahan, 2025), and it remains agnostic about whether such abilities imply anything like conscious experience (McClelland, 2024).
@@ 5.2 Numeric introspection as a complement to white-box methods
To make these contributions more useful for future research, we share an open-source library that allows this framework to be used with any concept, including methods for probe training, multi-probe recording, steering, and behavioral testing; we also propose a method for overcoming the natively collapsed rating outputs of models when asked to self-report emotive states, thereby making these reports more informative and improving the usefulness of introspection as a tool.
@@ 5.3 Introspection is concept-specific and only locally improvable in our experiments
The cross-concept steering results are therefore best read as an existence proof, not as evidence for a single, general “introspection knob”. We show that modest representational interventions can in some cases improve introspective fidelity in targeted concept pairs, which is already important: it demonstrates that basal introspective performance is not always maximal, and that some failure cases may be solvable by intervention rather than showing a complete absence of introspective capacity. But we did not search systematically for a universal steering direction that improves introspection across concepts. In pilot analyses we also examined additional interpretable directions, including truthfulness- and authenticity-style directions, but did not observe robust cross-concept introspection gains there (results not shown). The current evidence therefore points toward local, pair-specific improvement rather than a single globally tunable faculty. However, future work should still treat the search for this global faculty as an open problem.
@@ 5.5 Generalization to other models and limitations of this work
Even under that constraint, we find that introspective capacity generalizes to different model sizes and families. First, we find that introspective capacity grows with scale in two out of the four concepts we tested, reaching near-optimal introspection in the 8B LLaMA model (Fig. 5). Second, we replicate the core results outside the LLaMA family using the wellbeing concept in Qwen, while Gemma shows weaker but still detectable evidence. Therefore, we claim that introspective ability is not confined to a single architecture, at least for the wellbeing concept; we did not re-run the entire analysis pipeline for all concepts and models because of compute constraints. It is also important to note that we detected introspective ability in relatively small models. Much of the strongest prior evidence for mechanistically grounded introspection has involved larger or more heavily scaffolded systems, whereas our main experiments center on a 3B model and still recover causal and temporal structure (Li et al., 2025; Rivera and Africa, 2025; Lindsey, 2026).
@@
In our experiments, the emotive internal states that we study are partly determined by the previous context of the conversation. One could therefore argue that the models could be merely tracking and reporting features of the dialogue rather than performing true introspection, which requires privileged access. Our same-concept steering experiments show that self-reports change when the corresponding internal state changes, even when conversational context is the same. This shows that some degree of privileged access is necessary to explain the model’s output. However, part of the information the model is using to give self-reported ratings could be coming purely from conversational context. Prior work has shown that models can perform introspection in settings that do not depend on any specific natural conversation history (Lindsey, 2026; Pearson-Vogel et al., 2026; Li et al., 2025), which supports the claim that reports measured here are truly introspective. But it’s important to note that, to test the introspective capacity of models in the contexts where they will be used (e.g. during conversations and over time), it is necessary to study internal states that are related to that context, as we do in this work.